Find the length of a diagonal of a rectangle that is 16 units long and 12 units wide.
Answer by rwm(914)

This is classic Pythagorean theorem where the diagonal is the hypotenuse.
a^2+b^2=c^2
16^2+12^2=c^2
256+144=c^2
400=c^2
20=c
this is also the classic 3, 4,5 triangle multiplied by 4
12,16,20
Which you should know by heart